# T-FREE: Tokenizer-Free Generative LLMs via Sparse Representations for Memory-Efficient Embeddings

### v1

@@ Abstract
T-FREE: Tokenizer-Free Generative LLMs via
@@
Björn Deiseroth1,2,3  Manuel Brack2,4  Patrick Schramowski2,3,4 Kristian Kersting2,3,4Samuel Weinbach1 1 Aleph Alpha @ IPAI     2 Technical University Darmstadt 3 Hessian Center for Artificial Intelligence (hessian.AI) 4 German Center for Artificial Intelligence (DFKI)
@@
Figure 1: Method comparison of classic Tokenization (left) and T-Free (right) for text encoding (top) and decoding (bottom). Classic tokenizers learn a single-label vocabulary, i.e. a token is bijectively mapped into a single entry of the vocabulary. Instead, T-Free uses a bijective multi-label mapping over multiple activations of hashed character trigrams. As T-Free explicitly models morphological similarities, it enables compression of the embedding layer.
@@ 1 From Text Representations For Machine Learning
Although this approach has served the LLM community well, and influential characters target to tokenize all kinds of data to “lead a new industrial revolution” [^1] https://x.com/tsarnick/status/1801884651986030820?s=12&t=5I__mymj5rXz7lxfplR8Gg, it has significant inherent weaknesses. For one, tokenizers require dedicated training and, as such, additional computational resources. Design choices and errors at this stage can negatively impact the downstream model Ali et al. (2023). Any tokenizer’s vocabulary is heavily optimized for the reference corpus, leading to strong drops in performance for, e.g., underrepresented languages. We also show that the resulting vocabulary of tokenizers is poorly utilized, where up to 34% of tokens are near duplicates with limited additional information. Despite that, the corresponding embeddings are trained independently. These issues have caused a significant expansion in the size of vocabularies and their corresponding embedding layers, with billions of parameters being allocated exclusively for text encoding and decoding.
To remedy these issues and challenge the traditional views, we propose a paradigm shift on how we embed and decode text for LLMs. We present tokenizer-free sparse representations for memory-efficient embeddings (T-Free) as an alternative to tokenizers. We directly embed each word in the input text with sparse activation patterns over hashed character triplets. Consequently, we eliminate the need for subword tokens, thus retaining near-optimal performance across languages. Additionally, T-Free explicitly models character overlaps between morphologically similar words without the need to learn an embedding for each variant from scratch. We argue that the converged embedding of such similar words should remain close and, thus, can be heavily compressed [^2] The English language contains about $500k$ words, while “native fluency” is achieved at $10k$ words Nation (2006).. This exploitation of similarities allows T-Free to reduce the size of the embedding layers by $87.5\%$ [^3] Compared to our $64k$ unigram baseline. and the average encoding length of text by $56\%$ [^4] Compared to Mistral $32k$ avg. of EN, DE, RU, VI, AR.. In addition to the inherent benefits of T-Free, the approach remains highly competitive on standard downstream model performance benchmarks. Finally, for transfer learning to an unseen language, the T-Free model quickly improves performance, while the tokenizer baseline shows only minor adaptation.
@@ 3.2 Training Objective & Text Decoding
Analogously, for decoding the next token with T-Free, we first assemble a dictionary of all possible next words and pre-compute their activation patterns. Importantly, only $n\cdot m$ out of $v$ entries will be non-zero for each word, and since we choose $m<\!<v$, the dictionary matrix can be encoded as a sparse matrix, thus improving performance. We multiply this dictionary matrix with the predicted logits values of the LLM to finally obtain the argmax prediction of the next token.
3.3 Bridging the Gaps
@@ 4.1 Experimental Details
We built T-Free’s prediction dictrionary, from the top $80k$ words that occurring in slimpajama, and additional top $20k$ words from the German Occiglot data.
[…]
@@
4.3 T-Free has by design no duplicates
@@
4.4 T-Free has less fertility across, and is more adaptive to new languages
@@
Figure 3: Continual pre-training performance. Trained are $3B$ models on English slimpajama data for $90k$ steps (“baseline”), and continued on German occiglot data for $20k$ steps. Plotted are the average scores of two benchmarks available in German and English: Hellaswag and Arc-Challenge. Notably, T-Free outperforms in German already with the baseline. Within $20k$ continued steps, T-Free improves by $5\%$ on average in 0 and 2-shot, while the classic tokenizer approach barely improves. (Both models slightly drop performance in English, albeit the tokenizer version more drastically.)
@@ 6 Related Work
Few alternatives to BPE and Unigram have been found in recent LLMs and research. The naive approach of splitting the input text into bytes or characters maximizes fertility and thus increases computational requirements. Consequently, prior research has proposed methods for merging bytes, e.g., through state-space models Wang et al. (2024). However, these approaches still result in performance degradation. Finally, linguistically motivated approaches have built tokenizers based on known morphological rules Jabbar (2024). However, these methods are usually tailored to specific applications and are usually too costly and error-prone for large, general-purpose models.
Other works on weight tying, have halved the parameters of embedding and head layers by using the same matrix in both Press and Wolf (2017). Currently, LLMs do not apply weight tying, though, due to its negative impact on performance and the available compute.
@@ 7 Conclusion
In this work we present T-Free, an alternative to tokenizers with a simple and explicitly modeled robust hash function on words. It removes the need and pitfuls to limit “a models potential” to a “pre-pre-trained” tokenizer. We, moreover, fundamentally shift the established target of training language models, previously designed as a single-label problem, into a multi-label prediction based on word similarities. Similarities in particular include leading whitespaces and uppercase variations, for which tokenizers add specific tokens that are independently trained from scratch. These contributions allow us to train language models more robust, more adaptable when continuing pre-training with a new language, and with a significantly (to 12.5%) reduced parameter size without a decrease in benchmark scores. Due to the special role of the matrices, the latter in particular allows one to increase micro-batchsize, which further accelerates training time. Finally, the consequent convolution-like encoding achieves SOTA fertility scores across most languages and enables by design synergies to similar language groups. We demonstrated the latter showing that our $3B$ almost achieved “native-language” performance after a small amount of language-transfer training steps, in contrast to the tokenizer baseline.
@@
We expect T-Free to experience some numerical instabilities for very long words since single-word embeddings are calculated as the sum of their $n\cdot m$ activations. However, less than $2\%$ of the entire slimpajama dataset contains words with more than $10$ characters (cf. App. I), and we did not encounter any issues with the benchmarks. Consequently, such potential instabilities remain statistically insignificant. Nonetheless, we could adequately tackle long outliers with an additional split rule based on the words length.
Similarly, we did not thoroughly study the effect of repetitive trigrams in words. These did also not occur frequently enough to have any measurable effect on our experiments. As of now, we only accumulate a word pattern in a binary fashion, not accounting for trigrams appearing multiple times in a single word. As a fallback, one could again, split words at the position of repetitions.
Although T-Free’s fertility on code is on par with that of LLama2 (cf. App. E), it could be further improved by explicitly modeling code patterns. In this work, we have focused on natural language and leave detailed evaluations of T-Free in downstream coding tasks for future research. Furthermore, we did not investigate languages entirely relying on Unicode byte-encodings, such as Chinese.
@@ Appendix F Token Overlap/Duplicates
Ours 1.1636$\circ$ 1.1829 1.2363 1.1695 1.1274 1.3386 1.4001 1.0863$\bullet$
@@
Model Overlap (Total) (%) Capitalization (%) Whitespace (%) Digits (%)
Unigram Baseline (32k) 32.99 21.44 11.76 0.0
Unigram Baseline (64k) 35.24 23.27 13.47 0.0
BPE Baseline (32k) 32.12 21.30 13.85 0.0
BPE Baseline (64k) 35.32 23.82 15.52 0.0
@@
Phi-2: (50k) 23.23 12.91 16.89 3.32
DBRX (100k) 24.87 23.77 16.17 1.10
Gemma (256k) 34.68 20.27 20.50 0.04
Command-R (255k) 15.31 15.31 14.00 0.00
GPT-2 (50k) 25.25 21.93 16.99 3.32
Our 0 0 0 0
@@ Appendix H Hyperparameter Ablations
Albeit pretty scarse, some more hyper-parameter ablations are found in Fig. 6,7.
@@ Appendix I Some Statistics
We used the code of the eleuther eval harness, and evaluated each benchmark in 0-shot and 2-shot. All 18 benchmarks are found in Fig. 11 and Fig. 12.

### v2

# T-FREE: Subword Tokenizer-Free Generative LLMs via Sparse Representations for Memory-Efficient Embeddings

@@ Abstract
T-FREE: Subword Tokenizer-Free Generative LLMs via
@@
Björn Deiseroth1,2,3  Manuel Brack2,4  Patrick Schramowski2,3,4 Kristian Kersting2,3,4Samuel Weinbach1 1 Aleph Alpha @ IPAI     2 Technical University Darmstadt 3 Hessian Center for Artificial Intelligence (hessian.AI) 4 German Research Center for Artificial Intelligence (DFKI)
@@
Figure 1: Method comparison of classic Tokenization (left) and T-Free (right) for text encoding (top) and decoding (bottom). Classic subword tokenizers learn a single-label vocabulary, i.e. a token is bijectively mapped into a single entry of the vocabulary. Instead, T-Free uses a bijective multi-label mapping over multiple activations of hashed character trigrams. As T-Free explicitly models morphological similarities, it enables compression of the embedding layer.
@@ 1 From Text Representations For Machine Learning
Although this approach has served the LLM community well, and influential characters target to tokenize all kinds of data to “lead a new industrial revolution” [^1] https://x.com/tsarnick/status/1801884651986030820?s=12&t=5I__mymj5rXz7lxfplR8Gg, it has significant inherent weaknesses. For one, subword tokenizers require dedicated training and, as such, additional computational resources. Design choices and errors at this stage can negatively impact the downstream model Ali et al. (2023). Any tokenizer’s vocabulary is heavily optimized for the reference corpus, leading to strong drops in performance for, e.g., underrepresented languages. We also show that the resulting vocabulary of subword tokenizers is poorly utilized, where up to 34% of tokens are near duplicates with limited additional information. Despite that, the corresponding embeddings are trained independently. These issues have caused a significant expansion in the size of vocabularies and their corresponding embedding layers, with billions of parameters being allocated exclusively for text encoding and decoding.
To remedy these issues and challenge the traditional views, we propose a paradigm shift on how we embed and decode text for LLMs. We present tokenizer-free sparse representations for memory-efficient embeddings (T-Free) as an alternative to subword tokenizers. We directly embed each word in the input text with sparse activation patterns over hashed character triplets. Consequently, we eliminate the need for subword tokens, thus retaining near-optimal performance across languages. Additionally, T-Free explicitly models character overlaps between morphologically similar words without the need to learn an embedding for each variant from scratch through a one-to-one bijection. The backbone of the language model will remain free of subword tokenization as we directly encode the textual representation. We argue that the converged embedding of such similar words should remain close and, thus, can be heavily compressed [^2] The English language contains about $500k$ words, while “native fluency” is achieved at $10k$ words Nation (2006).. This exploitation of similarities allows T-Free to reduce the size of the embedding layers by $87.5\%$ [^3] Compared to our $64k$ unigram baseline. and the average encoding length of text by $56\%$ [^4] Compared to Mistral $32k$ avg. of EN, DE, RU, VI, AR.. In addition to the inherent benefits of T-Free, the approach remains highly competitive on standard downstream model performance benchmarks. Finally, for transfer learning to an unseen language, the T-Free model quickly improves performance, while the tokenizer baseline shows only minor adaptation.
@@ 3.1 Text Encoding
Note again, that we utilize a significantly smaller number of embeddings than there are trigrams. While their hashes may naturally overlap, we ensure the uniqueness of the entire patterns through the $m$ simultaneous hashes. As we ensure that trigram encodings do not collide, neither will the word encodings.
@@ 3.2 Training Objective & Text Decoding
Next token decoding is shown in Fig. 2. We first assemble a dictionary of all possible next words and pre-compute their activation patterns. Importantly, only $n\cdot m$ out of $v$ entries will be non-zero for each word, and since we choose $m<\!<v$, the dictionary matrix can be encoded as a sparse matrix, thus improving runtime. In addition, note the pattern similarity between similar words, as previously described. The last hidden layers’ output $h$ is sigmoided, and multiplied with the dictionary matrix. Finally, we compute the average sigmoid value per dictionary entry, $h^{\prime}$, to sample the next word, e.g. using standard argmax. Overall, for a dictionary with $512k$ entries, this procedure only marginally increases the decoding runtime due to the sparse property of the activation patterns. Further description, along with pseudocode, detailed depictions, and step-wise runtime analysis can be found in App. 5.
Note that the decode matrix is not required during training, and can dynamically be exchanged. We generate it by sampling the top-$500k$ occurring words in the training dataset, and dynamically adding the missing words of the prompt.
3.3 Distinctions of paradigm shift
@@
The predictable words are still derived from a dictionary. However, this vocabulary list is exchangeable, and is not required during training. As such, depending on the use-case, it may be kept in reasonable sizes. Moreover a hierarchical decoding exploiting morphological structures can straightforward be implemented, e.g. first decoding lowercase words, and then uppercase variations (or similarly grouping by stems or endings).
@@
Figure 2: Example of the next word prediction with T-Free. To the list of predictable words of dimension $d$ we generate once the corresponding patterns within the available vocabulary size $v$, as described in the encoding step $2$ of Sec. 3.1. Note how morphologically close words will generate overlapping patterns. The element-wise sigmoid values of the output of the last hidden layer, $\sigma(h)$, is multiplied with this pattern matrix using standard dot product. Finally, we use $h^{\prime}$ for the sampling process, the average sigmoid value of a word. C.f. App. A for further details.
@@ 4.1 Experimental Details
We built T-Free’s prediction dictionary, from the top $80k$ words that occurring in slimpajama, and additional top $20k$ words from the German Occiglot data.
@@
[…]
4.3 T-Free no duplicates by design
@@
4.4 T-Free has lower fertility across, and is more adaptive to new languages
@@
Figure 4: Continual pre-training performance. Trained are $3B$ models on English slimpajama data for $90k$ steps (“baseline”), and continued on German occiglot data for $20k$ steps. Plotted are the average scores of two benchmarks available in German and English: Hellaswag and Arc-Challenge. Notably, T-Free outperforms in German already with the baseline. Within $20k$ continued steps, T-Free improves by $5\%$ on average in 0 and 2-shot, while the classic tokenizer approach barely improves. Both models slightly drop performance in English, albeit the tokenizer version more drastically. Full evaluations are found in Appendix Tab. 8,9,10.
@@ 6 Related Work
Few alternatives to BPE and Unigram have been found in recent LLMs and research. Tay et al. (2022) propose a gradient-based trainable tokenization module in contrast the otherwise statistical based approach.
The naive approach of splitting the input text into bytes or characters maximizes fertility and thus increases computational requirements. Yu et al. (2023) employ a mix of multiple models to improve this drawback of byte-wise processing. I.p. they introduce a fixed character-embedding aggregation and a second character-decoder model. However, they use a fixed byte-width that is processed at once, which is not aligned with word splits.
Consequently, prior research has proposed methods for merging bytes, e.g., through state-space models Wang et al. (2024). However, these approaches still result in performance degradation. Finally, linguistically motivated approaches have built tokenizers based on known morphological rules Jabbar (2024). However, these methods are usually tailored to specific applications and are usually too costly and error-prone for large, general-purpose models.
Bojanowski et al. (2017) in particular discusses how adding subword informations, such as trigrams, enriches the encoding of words and leads to reliable compressions. Svenstrup et al. (2017) conduct research on the overloading of different hashfunctions to further improve and compress embedding representations. Xue and Aletras (2022) train BERT-style encoder models based on a different set of hashes on words. Clark et al. (2022) propose a multistage encoding scheme that uses hash functions and convolutions to enhance the BERT-encodings of words.
Another line of work to reduce the vocabulary parameter count is the utilization of weight tying, effectively halving it, as the embedding and head layers become “tied” to the same matrix Press and Wolf (2017). However, the effects on performance are still not sufficiently explored, and it arguably imposes a more difficult training objective.
@@ 7 Conclusion
In this work we present T-Free, an alternative to subword tokenizers with a simple and explicitly modeled robust hash function on words. It removes the need and pitfalls to limit “a models potential” to a “pre-pre-trained” vocabulary. We, moreover, fundamentally shift the established target of training language models, previously designed as a single-label problem, into a multi-label prediction based on word similarities. Similarities in particular include leading whitespaces and uppercase variations, for which subword tokenizers add specific tokens that are independently trained from scratch. These contributions allow us to train language models more robust, more adaptable when continuing pre-training with a new language, and with a significantly (to 12.5%) reduced parameter size without a decrease in benchmark scores. Due to the special role of the matrices, the latter in particular allows one to increase micro-batchsize, which further accelerates training time. Finally, the consequent convolution-like encoding achieves SOTA fertility scores across most languages and enables by design synergies to similar language groups. We demonstrated the latter showing that our $3B$ almost achieved “native-language” performance after a small amount of language-transfer training steps, in contrast to the tokenizer baseline.
@@
We expect T-Free to experience some numerical instabilities for very long words since single-word embeddings are calculated as the sum of their $n\cdot m$ activations. However, less than $2\%$ of the entire slimpajama dataset contains words with more than $10$ characters (cf. App. I), and we did not encounter any issues with the benchmarks. Consequently, such potential instabilities remain statistically insignificant. Nonetheless, we could adequately tackle long outliers with an additional split rule based on the words length or at the occurrence of repetitions. Subword tokenizers already demonstrate that such approaches will work, even when tokens are at first glance meaningless and underutilized—and again, these cases remain outliers. Moreover, a hybrid setup utilizing a large tokenizer (>512k tokens) with T-Free for optimized memory footprint is depicted in Figure 16.
Similarly, we did not thoroughly study the effect of repetitive trigrams in words. These did also not occur frequently enough to have any measurable effect on our experiments. As of now, we only accumulate a word pattern in a binary fashion, not accounting for trigrams appearing multiple times in a single word. As a fallback, one could again, split words at the position of repetitions. Another promising direction would overload embeddings with positional encodings similar to rotary Su et al. (2024).
Although T-Free’s fertility on code is on par with that of LLama2 (cf. App. E), it could be further improved by explicitly modeling code patterns. In this work, we have focused on natural language and leave detailed evaluations of T-Free in downstream coding tasks for future research. Furthermore, we did not investigate languages entirely relying on Unicode byte-encodings, such as Chinese. However, as they seemingly work out-of-the-box with subword tokenizers, we do not expect issues here by splitting them character/ word-wise. In particular for asian symbols, additionally translating the symbols to its romanization through the phonetic alphabet such as pinyin may further improve the synergies of word encodings.
@@ Appendix A T-Free Algorithm
Finally, refer to Figure. 14,15 for a step-wise comparison of the computation step, parameters and runtimes. Figure 16 shows a “hybrid” mode, in which embody a classical subword-tokenizer as a text preprocessing step, but utilize T-Free to keep the “tokenizer free LLM backbone”. Arguably, this approach benefits from a compressed embedding layer, and the tokenizer may easier be exchanged afterwards—the encoding of the text-chunks in the backbone will be kept as proposed.
@@ Appendix F Token Overlap/Duplicates
NeMo-Tekken (131k) 1.2313 1.5178 1.3061 1.3845 1.4171 2.0521 1.8378 1.6045
Ours 1.1636$\circ$ 1.1829 1.2363 1.1695 1.1274 1.3386 1.4001 1.0863
@@
Model/Tokenizer Portion of duplicate tokens (%) $\downarrow$
Total Cap. Space Digits
Unigram Baseline (32k) 32.99 21.44 11.76 0.00
Unigram Baseline (64k) 35.24 23.27 13.47 0.00
BPE Baseline (32k) 32.12 21.30 13.85 0.00
BPE Baseline (64k) 35.32 23.82 15.52 0.00
Phi-2: (50k) 23.23 12.91 16.89 3.32
DBRX (100k) 24.87 23.77 16.17 1.10
GPT-2 (50k) 25.25 21.93 16.99 3.32
Gemma (256k) 34.68 20.27 20.50 0.04
Command-R (255k) 15.31 15.31 14.00 0.00
@@
NeMo-Tekken (131k) 23.12 13.30 11.99 0.00
T-Free (Ours) 0 0 0 0
@@ Appendix H Hyperparameter Ablations
Albeit pretty scarse, some more hyper-parameter ablations are found in Fig. 7,8.
@@ Appendix I Some Statistics
We used the code of the eleuther eval harness, and evaluated each benchmark in 0-shot and 2-shot. All 18 benchmarks, namely arc (easy and challenge), hellaswag, winogrande, triviaqa, xnli, truthfulqa, boolq, copa, openbook, piqa, multirc, lambada (openai and standard), race, rte, wic, webqs are visualized in Fig. 12 and Fig. 13 for a baseline model trained on english slimpajama only and continued finetuning on german occiglot. Arc-challenge, hellaswag, xnli and truthfulqa are also evaluated in german translations. Detailed numbers can be found in Tab. 8,9 and 10.
@@
Model english benchmarks german benchmarks
$arc_{ch}$ $hella$ $xnli$ $tf_{mc1}$ $tf_{mc2}$ $arc_{ch}$ $hella$ $xnli$ $tf_{mc1}$ $tf_{mc2}$
base T-Free 34.5/36.2 63.1/62.2 33.4/34.8 25.9/- 44.1/- 24.3/25.0 27.7/28.1 37.1/32.7 31.1/- 44.3/-
Unigram 32.6/34.8 64.5/64.8 31.4/32.1 22.3/- 38.8/- 22.5/23.5 26.8/27.0 35.7/33.1 31.3/- 43.1/-
5k T-Free 34.7/35.1 63.0/61.9 32.5/33.2 24.7/- 44.1/- 30.9/32.5 28.3/28.4 38.1/34.9 31.1/- 44.3/-
Unigram 31.3/36.1 63.8/64.7 31.1/31.5 22.6/- 39.6/- 25.2/25.1 26.6/27.2 37.3/35.0 31.3/- 43.2/-
10k T-Free 34.1/35.1 61.9/61.1 32.2/33.8 24.2/- 44.1/- 32.1/33.4 29.6/29.7 35.7/34.0 30.0/- 44.3/-
Unigram 31.7/35.0 63.9/64.5 32.0/32.9 22.5/- 39.4/- 25.1/26.6 26.9/27.8 37.0/33.5 31.5/- 43.1/-
15k T-Free 33.9/36.3 61.8/61.8 32.3/34.9 24.2/- 44.1/- 33.2/33.9 30.7/30.4 35.9/35.1 30.8/- 44.3/-
Unigram 30.6/35.8 63.5/63.9 31.4/30.5 22.6/- 39.4/- 25.3/26.4 26.9/27.6 36.5/33.4 30.2/- 43.2/-
20k T-Free 35.3/35.5 62.2/59.9 32.4/35.1 24.4/- 44.1/- 32.0/33.7 31.9/31.5 35.6/35.9 29.1/- 44.2/-
Unigram 30.6/33.2 58.9/59.3 33.1/31.9 22.6/- 40.2/- 25.7/26.5 26.3/26.1 36.0/36.9 30.7/- 44.4/-
Table 8: Accuracy scores of english and german translated benchmarks for continued pre-training. First value denotes 0-shot, second value 2-shot (if available). Notably, the T-Free baseline model slightly outperforms (or performs on par with) the Unigram baseline model on all of these tasks. On german evals of arc and hellaswag, the T-Free baseline outperforms Unigram, and achieves larger gains during continued training on the german/ english data mix. The german versions of xnli and truthfulqa mostly remain unchainged.
Model english benchmarks
$arc_{ez}$ $boolq$ $copa$ $wino$ $obook$ $piqa$ $trivia$ $mrc$
base T-Free 62.1/69.4 62.9/59.8 70.0/71.0 58.7/58.3 35.4/35.0 74.0/73.8 15.7/29.3 36.0/38.1
Unigram 61.3/68.4 59.2/55.8 72.0/68.0 59.0/60.4 38.8/37.6 76.5/76.0 17.2/25.6 40.9/42.4
5k T-Free 60.8/64.8 62.6/60.7 70.0/70.0 59.0/57.9 35.2/35.4 73.0/73.0 13.1/22.9 37.8/36.9
Unigram 62.0/68.0 58.7/53.5 70.0/70.0 59.2/61.6 36.4/39.0 77.1/75.9 15.2/25.5 34.1/43.5
10k T-Free 58.8/66.3 60.5/60.6 71.0/73.0 59.2/57.8 34.2/34.2 74.4/72.9 10.9/22.8 38.0/36.6
Unigram 59.0/67.7 61.2/61.4 70.0/74.0 58.4/63.2 36.0/37.8 76.2/75.9 11.4/25.6 34.1/44.6
15k T-Free 60.2/66.6 63.8/60.0 72.0/76.0 58.6/59.3 34.4/35.6 73.9/73.3 13.8/24.9 38.6/37.3
Unigram 59.4/67.8 59.0/52.6 72.0/68.0 59.9/62.2 38.4/38.6 76.2/75.5 15.7/25.4 34.0/43.8
20k T-Free 62.5/66.0 62.0/62.6 70.0/72.0 57.9/57.1 35.6/36.5 74.6/74.1 12.5/22.4 39.7/36.7
Unigram 60.1/66.0 63.8/62.9 72.0/71.0 58.5/61.8 35.0/37.4 73.5/74.2 13.4/21.3 34.0/42.7
Table 9: Accuracy scores of english benchmarks for continued pre-training. First value denotes 0-shot, second value 2-shot. Notably, the T-Free model performs on par to the Unigram model on all of these tasks, throughout the entire continued training.
Model english benchmarks
$lbd^{oai}$ $lbd^{oai}_{clz}$ $lbd^{stdr}$ $lbd^{stdr}_{clz}$ $race$ $rte$ $wic$ $webqs$
base T-Free 53.9/46.4 18.5/42.9 48.0/44.8 12.2/39.3 37.8/39.1 58.5/55.2 50.0/53.8 5.7/26.5
Unigram 54.3/47.6 5.2/35.8 47.8/44.9 4.7/24.9 38.6/39.7 56.7/54.9 49.8/49.5 5.2/14.6
5k T-Free 52.7/45.7 6.6/41.9 45.6/43.0 7.8/44.0 38.4/40.8 59.6/55.6 50.0/54.5 7.6/20.8
Unigram 54.4/47.1 3.9/33.4 44.6/43.2 3.5/24.0 38.7/38.9 54.5/53.4 50.0/53.6 3.0/14.3
10k T-Free 53.1/47.3 13.1/42.8 46.2/44.7 7.5/44.7 37.3/39.9 56.7/54.9 50.0/54.7 5.4/19.8
Unigram 51.6/47.6 4.3/34.1 46.0/44.3 5.2/29.4 39.0/39.9 54.5/54.5 49.7/53.9 3.0/13.8
15k T-Free 52.8/46.2 12.5/40.4 44.7/42.6 9.5/42.7 37.1/40.0 57.4/50.9 50.0/54.2 7.6/21.8
Unigram 53.6/47.7 4.8/30.1 46.5/42.0 5.4/23.5 39.0/38.5 53.4/53.4 49.8/50.6 3.0/14.6
20k T-Free 53.5/46.0 9.9/36.2 46.1/44.6 11.9/41.6 37.4/39.1 54.9/55.6 50.0/54.5 7.4/20.3
Unigram 52.2/45.2 3.7/30.8 40.1/36.4 1.0/16.6 38.8/38.8 54.2/57.0 49.5/53.9 3.2/13.1
Table 10: Accuracy scores of english benchmarks for continued pre-training. First value denotes 0-shot, second value 2-shot. The T-Free model performs on par with the Unigram model on all of these tasks, throughout the entire continued training. Notably, the clozed variants of lambada are most fragile, at which T-Free outperforms.
[…]
Figure 15: “Greedy” text-decoding example for T-Free (top) and classic decoder LLMs (bottom). T-Free applies a head of significantly reduced parameters which results in less dense matrix multiplications and smaller vector sizes. As an additional step, during inference, T-Free computes the average activation score $h^{\prime}_{out}$, which is sparsely computed by multiplying (and averaging) the once precomputed decodable dictionary with the sigmoid scores of the head. Finally, in both cases argmax is taken to lookup the resulting word.
Figure 16: Hybrid “T-Free” (tokenizer-free/adaptable) LLM Backbone applying large scale (500k+) tokenizers. Major advantages of a T-Free backbone in a hybrid setting are the compression of embedding and head matrices, and the potential flexibility to lateron exchange (with some finetuning) the tokenizer — the backbone remains with the same tokenizer-free encoding rules.